∠A and ∠B are complementary. Find m∠A and m∠B.
M∠A= (3x+2)
M∠B= (x-4’)
M∠A=
M∠B=

∠A and ∠B are supplementary. Find m ∠A and m∠B. M∠A=(8x+100) m∠B(2x+50)

M ∠ A =
M ∠B=
1
Expert's answer
2020-09-15T16:27:59-0400

Given

"m\u2220A=(3x+2)"

"m\u2220B=(x-4)"

And

∠A and ∠B are complementary angel.


So, we know that

"m\u2220A+m\u2220B=90\u00b0"


After the plug in the value

m∠A and m∠B.

"(3x+2)+(x-4)=90\u00b0"

Combining the like tems

"4x-2=90\u00b0"

"4x=90\u00b0"

"x=23\u00b0"

Now we substitute the value of x in the ∠A and ∠B.

"m\u2220A=(3\u00d723+2)=71\u00b0"

"m\u2220B=23-4=19\u00b0"



Given

"m\u2220A=(8x+100)"

"m\u2220B=(2x+50)"

And


∠A and ∠B are supplementary angel.

"\u2220A+\u2220B=180\u00b0"

After plug in the value of m∠A and m∠B.

"(8x+100)+(2x+50)=180\u00b0"

Combining the like tems

"10x+150=180\u00b0"

"10x=30\u00b0"

"x=3\u00b0"

Now we substitute the value of x in the ∠A and ∠B.

"m\u2220A=8\u00d73+100=124\u00b0"

"m\u2220B=2\u00d73+50=56\u00b0"